OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 4x - 16 over 12 - 3x?

OpenStudy (shubhamsrg):

x-4 gets cancelled if you take 4 from above and 3 from denom.

OpenStudy (anonymous):

4x-16=4(x-4) and 12-3x=-3(x-4) .. and cancel x-4

OpenStudy (shubhamsrg):

infact,,either -4 from num or -3 from denom..

OpenStudy (anonymous):

That's what I thought. Thank you. :)

OpenStudy (shubhamsrg):

nevermind.. ;)